Operations Report – March 30th

Hi CapCom,

this is the Engineer Report on sol 6.

Best,
Miho XO crew191

Crew 191 Operations Report 30/03/2018

SOL:6

Name of person filing report: Wataru Okamoto
Non-nominal systems: Solar System
Notes on non-nominal systems: Solar system seems unusual from yesterday. P05 error code are there.

Generator (hours run):hours
Solar SOC 99% (Before generator is run at night)
Diesel under 25%
Propane 70psi.
Ethanol Free Gasoline (5 Gallon containers for ATV) ? 6gallons.
Water (trailer) 60% Water (static) 70%
Trailer to Static Pump used ? no
Water (loft) Static to Loft Pump used yes
Water Meter: 132861.4
Toilet tank emptied: yes
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –
Oil Added –
ATV Fuel Used: N/A
Hours the ATVs were Used today: N/A
Notes on ATVs: N/A
Deimos rover used: –
Hours:116.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: 100%
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: –
Hours: 27.9h
Beginning charge:100%
Ending charge: 100%

Currently charging: 100%
Opportunity rover used: Murakami, Okamoto
While driving became overheat twice.
Hours: 28.4h
Beginning charge: 100%
Ending charge: 86%
Currently charging: yes
Curiosity rover used: Venza, Kawamura
Hours:23.7h
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes
HabCar used and why, where? –
General notes and comments: Solar system seems unusual from yesterday. We had precise monitoring and also discussion about this problem with Dr. Shannon this morning. The findings of the monitoring are that solar charger system looks fine, but the number of monitor telling us wrong result for some a while. For example, 11:11 SOC has changes from 63% to 100% suddenly. We will keep monitoring.
Summary of internet: Under control
Summary of suits and radios: No problem
Summary of Hab operations: Shower heater doesn’t work.
Summary of GreenHab operations: No problem
Summary of ScienceDome operations: Only use checking the solar system.
Summary of RAMM operations: N/A
Summary of health and safety issues: No problem
Questions, concerns and requests to Mission Support:
Opportunity rover became overheat twice while driving. At that same time curiousity was fine. Is there any advice?

Operations Report March 29th

Crew 191 Operations Report 29/03/2018

SOL:5

Name of person filing report: Wataru Okamoto

Non-nominal systems: Solar System

Notes on non-nominal systems: Solar system seems unusual in this morning. Now it looks stable.

Generator (hours run):13hours

Solar SOC 93% (Before generator is run at night)

Diesel under 25%

Propane 72psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 7gallons.

Water (trailer) 60%

Water (static) 70%

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 16:00 132835.5 

Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: N/A

Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

Deimos rover used: –

Hours:116.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: 100%

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: –

Hours:27.9h

Beginning charge:100%

Ending charge: 100%

Currently charging: 100%

Opportunity rover used: Morisawa, Takeda

Hours: 27.4h

Beginning charge: 100%

Ending charge: 99%

Currently charging: 99%

Curiosity rover used: Tsukishiro

Hours:22.7h

Beginning charge: 100%

Ending charge: 97%

Currently charging: 97%

HabCar used and why, where? –

General notes and comments: Solar system seems unusual in this morning. We have discussion about this problem with Dr. Shannon and now it seems stable. We will keep monitoring.

Summary of internet: Under control

Summary of suits and radios: The surface of helmet polished by Plexus & microfiber clothes.

Summary of Hab operations: Shower heater doesn’t work.

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:

The PLEXUS splay is almost empty. We need another supply.

We will put some garbage bags into the engineering airlock tonight. Please take them out.

Operations Report – March 28th

Crew 191 Operations Report 28/03/2018

SOL:4

Name of person filing report: Wataru Okamoto

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Generator (hours run):12.6hours

Solar SOC % (Before generator is run at night)

Diesel under 25%

Propane 72psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 7gallons.

Water (trailer) 50%

Water (static) FULL

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 7:30 132802.5 

Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: N/A

Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

Deimos rover used: –

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: –

Hours:27.9h

Beginning charge:100%

Ending charge: 100%

Currently charging: 100%

Opportunity rover used: Tsukishiro, Takeda

Hours: 26.9h

Beginning charge: 100%

Ending charge: 100%

Currently charging: 100%

Curiosity rover used: Morisawa

Hours:22.2h

Beginning charge: 100%

Ending charge: 100%

Currently charging: 100%

HabCar used and why, where? –

General notes and comments: –

Summary of internet: Under control

Summary of suits and radios: One of radios’s (Midland GXT X-TRA TALK) internal microphone doesn’t work. But if connect an external microphone it can still be use.

Summary of Hab operations: Shower heater doesn’t work.

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:

– Please charge the Diesel gas.

– Please give us information how to repair the shower heater.

– We repair No.5 Backpack yesterday. Now We can use normaly.

– Please give us information of radio how to fix.

Operations Report – March 27th

Crew 191 Operations Report 27/03/2018

SOL:3

Name of person filing report: Wataru Okamoto

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Generator (hours run):13hours

Solar SOC 96% (Before generator is run at night)

Diesel 25%

Propane 70psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 7gallons.

Water (trailer) 60%

Water (static) 80%

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 7:50 6 132778.1->pump 14 16:00 10 132790.2

Toilet tank emptied: yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: – 350.2 45%->75%

Hours the ATVs were Used today: –

Notes on ATVs:-

Deimos rover used: –
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: –
Hours:27.9h
Beginning charge:100%
Ending charge: 100%
Currently charging: 100%

Opportunity rover used: –
Hours: 26.9h
Beginning charge: 100%
Ending charge: 100%
Currently charging: 100%

Curiosity rover used: –
Hours:22.1h
Beginning charge: 100%
Ending charge: 100%
Currently charging: 100%

HabCar used and why, where? –

General notes and comments: -I didn’t turn off the Diesel generator in this morning. I turned off the Diesel generator after 50 minuites. It’s my mistake.I’ll turn off when amount of battery is over 80%.

Summary of internet: Under control

Summary of suits and radios:No.5 Backpack Battery was 6.8V after charging. cSo another 13V battery was installed.2 Backpack belt have repaired

Summary of Hab operations:No problem

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:Please charge the Diesel gas.

Operations Report – March 22nd

SOL: 11

Name of person filing report: Bastien BAIX
Non-nominal systems: noisy pump, see end of report below
Notes on non-nominal systems: –

Generator (hours run): Turned off at 12.40 pm and turned on at 6.30 pm Solar – SOC 80% (Before generator is run at night)
Diesel – ~25%
Propane – 80%
Ethanol Free Gasoline (5 Gallon containers for ATV) – 1 ½ full jerrycans.

Water (trailer) – 0 gallons.
Water (static) – ~500 gallons
Water (third tank) – 0 gallons
Trailer to Static Pump used – NO
Water (loft) – Static to Loft Pump used – YES
Water Meter: 132 585.7
Toilet tank emptied: NO

ATV’s Used: –
Oil Added: NO
ATV Fuel Used: NO
Hours the ATVs were used today: _
Notes on ATVs: –

Deimos rover used: NO
Hours: 116.7
Beginning charge: 100%
Ending charge: –
Currently charging: YES

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: NO
Hours: 26.9
Beginning charge: 100%
Ending charge: –
Currently charging: NO

Opportunity rover used: YES
Hours: 26
Beginning charge: 100%
Ending charge: 93%
Currently charging: NO

Curiosity rover used: YES
Hours: 20.8
Beginning charge: 100%
Ending charge: 91%
Currently charging: YES

HabCar used and why, where: –
General notes and comments: –
Summary of internet: ~ 120 Mb remaining
Summary of suits and radios: –
Summary of Hab operations: –
Summary of GreenHab operations: –
Summary of ScienceDome operations: –
Summary of RAMM operations: –
Questions, concerns and requests to Mission Support:

Update on the pump: I made pictures of the pump (and the old pump too). – Model: Pentair Shurflo
– 115 VAC
– 60Hz
– 3,3 GPM /12,5 LPM of max flow
– 104 Watts
– Pressure: 45 PSI / 3,1 bar

Operations Report – 21st March

Crew 190 Operations Report 21th March 2018

SOL: 10

Name of person filing report: Bastien BAIX

Non-nominal systems: Two water pumps: tank-to-tank pump, interior Hab pump

Notes on non-nominal systems: tank-to-tank pump is taking significantly longer than normal (pump is used to transfer water from the exterior water tank to the interior tank). Interior pump, used to inject water into the Hab water system is making strange noises. Cause is unknown. Should we worry about it ?

Generator (hours run): Turned off at 10.40 am and turned on at … pm (not yet turned on).

Solar – SOC …% (Before generator is run at night)

Diesel – ~35%

Propane – 80%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1 ½ full jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~525 gallons

Water (third tank) – 0 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 529.4

Toilet tank emptied: YES

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.7

Beginning charge: 100%

Ending charge: –

Currently charging: YES

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: YES

Hours: 26.9

Beginning charge: 100%

Ending charge: 95%

Currently charging: YES

Opportunity rover used: YES

Hours: 25,7

Beginning charge: 100%

Ending charge: 90%

Currently charging: YES

Curiosity rover used: NO

Hours: 20.4

Beginning charge: 100%

Ending charge: –

Currently charging: NO

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 180 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Bastien and Martin working on their experiments

Summary of GreenHab operations: Mario and Fred working on their experiments

Summary of ScienceDome operations: Ariane, Michael, Maximilien and Fred working on their experiments

Summary of RAMM operations: Sophie working on her muon detector

Questions, concerns and requests to Mission Support: –

Operations Report – March 20th

Crew 190 Operations Report 20th March 2018

SOL: 9

Name of person filing report: Bastien BAIX

Non-nominal systems: –

Notes on non-nominal systems: –

Generator (hours run): Turned off at 9.37 am and turned on at 6.45 pm.

Solar – SOC 77% (Before generator is run at night)

Diesel – ~45%

Propane – 81%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1 ½ full jerrycans.

Water (trailer) – 0 gallons.

Water (static) – 550 gallons

Water (third tank) – 0 gallons

Trailer to Static Pump used – YES

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 470.8

Toilet tank emptied: YES

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: YES

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.6

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Notes on Deimos rover: the wheel is under reparation

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours: 26.2

Beginning charge: 71%

Ending charge: –

Currently charging: NO

Opportunity rover used: NO

Hours: 25

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Curiosity rover used: NO

Hours: 20.7

Beginning charge: 82%

Ending charge: –

Currently charging: YES

HabCar used and why, where: –

General notes and comments: The water level of the aquapony system is lowering due to evaporation, should we refill it ?

Summary of internet: ~ 0 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Power charging to rovers resolved. Fuses were being blown when two outlets were being used. With one cord the issue is resolved.

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Questions, concerns and requests to Mission Support: –

Operations Report – March 18th

Crew 190 Operations Report 18th March 2018

SOL: 7

Name of person filing report: Bastien BAIX

Non-nominal systems: Ventilation cover fell.

Notes on non-nominal systems:

Due to the extremely hard winds of the last days, the cover of a ventilation opening above the kitchen fell. We found it on the ground near the GreenHab. As you can see on our self-made 3D model enclosed to this mail, it is high. We don’t have a ladder long enough to fix it.

Generator (hours run): Turned off at 9.45 am (and turned on at 7:10pm).

Solar – SOC 85 % (Before generator is run at night)

Diesel – ~45%

Propane – 81%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~ 250 gallons

Water (third tank) – 500 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 355.8

Toilet tank emptied: YES

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.6

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Notes on Deimos rover: flat tire, not yet repaired

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: YES

Hours: 26.1

Beginning charge: 100%

Ending charge: 43%

Currently charging: NO

Opportunity rover used: YES

Hours: 25

Beginning charge: 100%

Ending charge: 30%

Currently charging: YES

Curiosity rover used: YES

Hours: 20.7

Beginning charge: 100%

Ending charge: 55%

Currently charging: YES

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 270 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: day off

Summary of GreenHab operations: –

Summary of ScienceDome operations: Fred working on their experiments.

Summary of RAMM operations: the muon detector of Sophie is remotely controlled from the Hab

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Operations Report – March 17th

Crew 190 Operations Report 17th March 2018

SOL: 6

Name of person filing report: Bastien BAIX

Non-nominal systems: –

Notes on non-nominal systems: –

Generator (hours run): Turned off at 9.36 am and turned on at 6.10 pm.

Solar – SOC 99% (Before generator is run at night)

Diesel – ~45%

Propane – 81%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~ 250 gallons

Water (third tank) – 500 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 303.3

Toilet tank emptied: NO

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.6

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Notes on Deimos rover: flat tire, not yet repaired

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: YES

Hours: 25.1

Beginning charge: 89%

Ending charge: 100%

Currently charging: YES

Opportunity rover used: NO

Hours: 24.1

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Curiosity rover used: YES

Hours: 19.3

Beginning charge: 86%

Ending charge: 100%

Currently charging: YES

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 250 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Ariane, Sophie & Bastien working on their experiments

Summary of GreenHab operations: –

Summary of ScienceDome operations: Martin, Maximilien & Fred working on their experiments.

Summary of RAMM operations: the muon detector of Sophie is remotely controlled from the Hab

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Operations Report – March 16th

Hello CapCom,

Here is my sixth engineering report, feel free to give me feedback:

Crew 190 Operations Report 16th March 2018

SOL: 5

Name of person filing report: Bastien BAIX

Non-nominal systems: –

Notes on non-nominal systems: –

Generator (hours run): Turned off at 8.30 am and turned on at 6.45 pm.

Solar – SOC 95% (Before generator is run at night)

Diesel – ~45%

Propane – 81%

Ethanol Free Gasoline (5 Gallon containers for ATV) – 2 empty jerrycans.

Water (trailer) – 0 gallons.

Water (static) – ~ 250 gallons

Water (third tank) – 500 gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 132 251.9

Toilet tank emptied: YES

ATV’s Used: –

Oil Added: NO

ATV Fuel Used: NO

Hours the ATVs were used today: _

Notes on ATVs: –

Deimos rover used: NO

Hours: 116.6

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Notes on Deimos rover: flat tire, not yet repaired

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: NO

Hours: 24.8

Beginning charge: 100%

Ending charge: –

Currently charging: NO

Opportunity rover used: NO

Hours: 24.1

Beginning charge: 100%

Ending charge: –

Currently charging: YES

Curiosity rover used: NO

Hours: 18.1

Beginning charge: 100%

Ending charge: –

Currently charging: YES

HabCar used and why, where: –

General notes and comments: –

Summary of internet: ~ 200 Mb remaining

Summary of suits and radios: –

Summary of Hab operations: Ariane, Mario & Bastien working on their experiments

Summary of GreenHab operations: –

Summary of ScienceDome operations: Martin, Maximilien & Fred working on their experiments.

Summary of RAMM operations: Sophie adjusting her muon detector

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Enjoy your evening,

Cheers,

Bastien Baix

Crew 190 Engineer – UCL to Mars 2018